OpenStudy (anonymous):

"the difference of two positive numbers is one-third of their sum. the sum of the squares of the two numbers is 180

OpenStudy (anonymous):

\[x - y = \left(\begin{matrix}1 \\ 3\end{matrix}\right) (x+y)\]